# Bounded geometry for PCF-special subvarieties

### v2

@@ 2.4. The critical height
defines a function on $\mathcal{M}_{d}^{cm}$, which Silverman called the critical height and conjectured that it is comparable with a Weil height on $\mathrm{M}_{d}\setminus\mathcal{L}_{d}$; see, for example, [Si2]. This was proved by Ingram in [In].
@@
Yuan and Zhang use their general equidistribution theorem [YZ, Theorem 5.4.3] to give an alternative proof of the arithmetic equidistribution of the points in $\mathrm{PCF}$ to the bifurcation measure $T_{\mathrm{bif}}^{\wedge 2d-2}$ in [YZ, §6.3], a result which was first obtained by Gauthier in [Ga3] (and was proved for the polynomial moduli spaces in [FG1]).
@@ 2.5. The Chow variety and maximal variation
Suppose that $\pi:\mathcal{X}\to V$ is a family of $r$-dimensional varieties in $Y$, parametrized by $\lambda$ in an irreducible quasi-projective variety $V$ of dimension $\ell\geq 1$. Assume that the generic fiber of $\mathcal{X}\to V$ is geometrically irreducible. We say that $\mathcal{X}\to V$ is maximally varying if the natural map from $V$ to the Chow variety $\mathrm{Ch}(Y,r,D)$ has finite fibers, where we send each irreducible fiber $X_{\lambda}:=\pi^{-1}(\lambda)$ to the point of $\mathrm{Ch}(Y,r,D)$ representing the cycle $[\overline{X}_{\lambda}]$ on $\overline{Y}$.
@@
Now suppose $\ell>1$ and that we know the result for maximally varying families with base dimension $<\ell$. Note that $\dim\mathcal{X}=\ell+r$ and $\dim\mathcal{X}_{V}^{\ell}=\ell(r+1)$. Choose any smooth point $x_{0}$ of $\rho_{1}(\mathcal{X})$ which is also smooth for a subvariety $X_{\lambda_{0}}$ it lies in, and which is regular for the maps $\rho_{1}:\mathcal{X}\to\rho_{1}(\mathcal{X})$ and $\pi_{1}:\rho_{\ell}(\mathcal{X}_{V}^{\ell})\to\rho_{1}(\mathcal{X})$ where the latter is the restriction of the projection $Y^{\ell}\to Y$ to the first coordinate. Here, ‘regular for the maps’ means that the fiber over $x_{0}$ has the expected dimension and there is a point in the fiber over $x_{0}$ mapping surjectively on tangent spaces to that of $x_{0}$.
@@ 3. Bifurcation measures on $\mathcal{X}_{V}^{\ell}$ and $\mathcal{X}_{V}^{\ell+1}$
In this section we construct nontrivial measures on $\mathcal{X}_{V}^{\ell}$ and $\mathcal{X}_{V}^{\ell+1}$ that are pullbacks of powers of the bifurcation currents in $\mathcal{M}_{d}^{cm}$, defined in §2.3. The measure on $\mathcal{X}_{V}^{\ell}$ will be defined in terms of $T_{\mathrm{bif}}$, while two measures on $\mathcal{X}_{V}^{\ell+1}$ will additionally involve choices of $r$-tuples of marked critical points. The proofs build on the strong rigidity result that the power $T_{\mathrm{bif}}^{\wedge r}$ of the bifurcation current is nonzero on every $r$-dimensional subvariety of $\mathrm{M}_{d}\setminus\mathcal{L}_{d}$, where $\mathcal{L}_{d}$ is the flexible Lattès locus; this is proved in [GOV, Lemma 6.8] for all $r$. (See the discussion in §2.3.)
We also appeal to descriptions of the support of powers $T_{\mathrm{bif}}^{\wedge r}$ in $\mathrm{M}_{d}$ in terms of critical orbit behavior obtained by Buff-Epstein in [BE] and Dujardin in [Du]. Precisely, suppose that $f_{1},\ldots,f_{m}$ is a collection of holomorphic families of maps on $\mathbb{P}^{1}$ parameterized by $\lambda$ in a complex manifold $\Lambda$, and let $\Phi:=(f_{1},\ldots,f_{m}):\Lambda\times(\mathbb{P}^{1})^{m}\to\Lambda\times(\mathbb{P}^{1})^{m}$ be the induced holomorphic map. We say that an $m$-tuple of marked points ${\bf a}=(a_{1},\ldots,a_{m}):\Lambda\to(\mathbb{P}^{1})^{m}$ is transversely (respectively, properly) prerepelling at $\lambda_{0}$ over $\Lambda$ if a $\Phi$-iterate of the graph of ${\bf a}$ intersects the graph of a repelling periodic point for $\Phi$ transversely (resp., properly) in $\Lambda\times(\mathbb{P}^{1})^{m}$ over $\lambda_{0}$.
When the marked points are critical points, with $f:=f_{1}=\cdots=f_{m}$, Dujardin proved that there is a dense set of parameters in $\operatorname{supp}T_{1}\wedge\cdots\wedge T_{r}$ in $\mathcal{M}_{d}^{cm}$ at which the critical point ${\bf c}=(c_{1},\ldots,c_{r})$ for $(f,\ldots,f)$ is transversely prerepelling over $\mathcal{M}_{d}^{cm}$ [Du, Theorem 1.1]. And conversely, if the $r$-tuple of critical points is properly prerepelling for $(f,\ldots,f)$ at some parameter over a parameter space of dimension $\geq r$, then $T_{1}\wedge\cdots\wedge T_{r}$ must be nonzero in that parameter space, as first observed in [BE] (where it was proved in the case $r=2d-2$ and assuming transversely prerepelling) and revisited in [Ga1, Theorem 6.2] and [BBD, Proposition 3.7] in more general contexts and without assuming transversality. See [DM1, Proposition 4.8] for the version we use.
@@ 3.1. Measure on $\mathcal{X}_{V}^{\ell}$
so that $\mu_{\ell}\not=0$.
@@ 3.2. Measures on $\mathcal{X}_{V}^{\ell+1}$
be as in Proposition 3.1 parameterized by $\mathcal{X}_{V}^{\ell}$, and let
@@
be the family of maps on $(\mathbb{P}^{1})^{\ell(r+1)+r}$ parameterized by $\mathcal{X}_{V}^{\ell+1}$, where $f_{1},\ldots,f_{\ell}$ are each repeated $r+1$ times, but the final coordinate $f_{\ell+1}$ is only repeated $r$ times.
@@
    the $r$-tuple $(c_{\ell,1},\ldots,\widehat{c_{\ell,i}},\ldots c_{\ell,r+1})$ is transversely prerepelling for $(f_{\ell},\ldots,f_{\ell})$ at $t_{\ell}$ over $X_{\lambda_{0}}$ and at $s_{\ell}$ over $X_{\lambda_{1}}$;
@@
    the $r$-tuple $(c_{\ell,1},\ldots,\widehat{c_{\ell,j}},\ldots c_{\ell,r+1})$ is properly prerepelling for $(f_{\ell},\ldots,f_{\ell})$ at $t_{\ell}$ over $X_{\lambda_{0}}$; and
@@
    $c_{\ell,i}$ is attracted to a parabolic periodic point for $f_{\ell}$ at $s_{\ell}\in X_{\lambda_{1}}$,
where $\widehat{c_{\ell,i}}$ means that coordinate is omitted. Let $\Pi_{1}:\mathcal{X}^{\ell}_{V}\times_{V}\mathcal{X}\to\mathcal{X}_{V}^{\ell}$ and $\Pi_{2}:\mathcal{X}_{V}^{\ell}\times_{V}\mathcal{X}\to\mathcal{X}$. Then
@@
Over $\lambda_{0}$, by the transversality, we know that $r$ of the remaining critical points, say $c_{\ell,1},\ldots,c_{\ell,r}$, are transversely prerepelling at $t_{\ell}$ for $f_{\ell}$ along $X_{\lambda_{0}}$.
We now aim to show that by modifying the parameter $x_{0}$ if necessary, we can arrange so that another $r$-tuple, say $c_{\ell,2},\ldots,c_{\ell,r+1}$, is properly prerepelling at $t_{\ell}$ for $f_{\ell}$ along $X_{\lambda_{0}}$. Note that transversality of the intersection $S_{1,1}\cap\cdots\cap S_{\ell-1,r+1}$ at the original $x_{0}$ persists in a neighborhood, so the first $(\ell-1)(r+1)$ critical points remain transversely prerepelling in a neighborhood of $x_{0}$ for $(f_{1},\ldots,f_{\ell-1})$.
Now consider the intersection
@@
We conclude that there is a parameter $\lambda^{\prime}$ near $\lambda_{0}$ so that critical points $c_{\ell,1},\ldots,c_{\ell,r}$ are transversely prerepelling at $x^{\prime}$ along $X_{\lambda^{\prime}}$ while $c_{\ell,2},\ldots,c_{\ell,r+1}$ are properly prerepelling at $x^{\prime}$ along $X_{\lambda^{\prime}}$. This $x^{\prime}$ defines the new $x_{0}$ for the statements (1) and (2) of the proposition.
As $c_{\ell,r+1}$ is bifurcating along the curve we called $\Gamma$, we know that there are baby Mandelbrot (or Multibrot) sets everywhere dense in its bifurcation locus (see for example [Mc2]); therefore, there is a nearby parameter $\lambda_{1}$ in the support of $\pi_{*}\mu_{\ell}$, where $\pi$ is the projection to $V$, so that $c_{\ell,r+1}$ is attracted to a parabolic cycle at $x_{1}:=\Gamma\cap X_{\lambda_{1}}$, while $c_{\ell,1},\ldots,c_{\ell,r}$ remain transversely prerepelling at $x_{1}$ along $X_{\lambda_{1}}$. This completes the proof of properties (1), (2), and (3).
To prove the non-vanishing of $\mu_{\ell}^{i}$ and $\mu_{\ell}^{j}$, consider the map $\Phi_{\ell+1}=f_{1}^{\times(r+1)}\times\cdots\times f_{\ell}^{\times(r+1)}\times f_{\ell+1}^{\times r}$ on $(\mathbb{P}^{1})^{\ell(r+1)+r}$ over $\mathcal{X}_{V}^{\ell+1}$. The argument above shows that we can label critical points of $f_{1},\ldots,f_{\ell+1}$ so that the graphs of the marked points
@@
Suppose that $\mathcal{X}\to V$ is a maximally varying family of $r$-dimensional varieties $X_{\lambda}\subset M_{d}$ parametrized by $V$ of dimension $\ell\geq 1$, defined over $\overline{\mathbb{Q}}$. As in §2.1, we may pass to a finite branched cover $\mathcal{M}_{d}^{cm}\to\mathrm{M}_{d}$ over which we can define a family of maps $f:\mathbb{P}^{1}\to\mathbb{P}^{1}$ with $22d-$ 2 marked critical points. Pulling back each $X_{\lambda}$ and possibly replacing $V$ with a finite branched cover, we will continue to use the notation $\mathcal{X}\to V$ for the family of varieties in $\mathcal{M}_{d}^{cm}$. Therefore, we may assume that we can define the family of product maps
@@
In what follows it is convenient to work with the Deligne pairing (of relative dimension dimension 00), of the invariant metrized line bundles restricted to graphs of critical marked points (see [YZ, §4.2] for metrics of the Deligne pairing and [YZ, §6.2.2] for the ‘restriction’ of the invariant metrized line bundle to a subvariety). By [YZ, Theorem 4.2.3], to each marked tuple of critical points ${\bf c}:\mathcal{X}^{\ell}_{V}\to(\mathbb{P}^{1})^{\dim\mathcal{X}^{\ell}_{V}}$ we can associate a metrized line bundle $\langle\overline{L}_{\Phi}|_{\Gamma_{\bf c}}\rangle\in\widehat{\mathrm{Pic}}(\mathcal{X}^{\ell}_{V})_{\mathbb{Q},\mathrm{nef}}$, where $\Gamma_{\bf c}$ denotes the graph of ${\bf c}$ in $\mathcal{X}^{\ell}_{V}\times(\mathbb{P}^{1})^{\dim\mathcal{X}^{\ell}_{V}}$. Moreover, we let
@@
on $\mathcal{X}^{\ell+1}_{V}$ for $i=1,2$. By [YZ, Lemma 6.2.1, Theorem 4.1.3], the heights associated to $\overline{M}_{i}$ detect PCF points; see also [YZ, Lemma 6.3.2]. More precisely, for a point $(\lambda_{0},t_{1},\ldots,t_{\ell+1})\in\mathcal{X}^{\ell+1}_{V}(\overline{\mathbb{Q}})$ we have
@@
[…]
@@
as measures on $X_{\lambda}$, for all $\lambda$ in the support of $\pi_{*}\mu_{\ell}$ in $V$. Indeed, suppose there exists $b\in V(\mathbb{C})$ where (4.8) doesn’t hold, and so that $\pi_{*}\mu_{\ell}(U)>0$ for every open neighborhood $U$ of $b$. Then we can find a continuous, non-negative, compactly supported function $\psi$ on $\mathcal{X}$ such that
[…]
@@
Recall from Proposition 3.2 that we have parameter $x_{1}=(\lambda_{1},s_{1},\ldots,s_{\ell})\in\operatorname{supp}\mu_{\ell}$ so that the tuple of marked critical points $(c_{1},\ldots,c_{r})$ is transversely prerepelling for $(f_{\ell},\ldots,f_{\ell})$ at $s_{\ell}$ along $X_{\lambda_{1}}$. This implies that $\mu_{\lambda_{1}}^{r+1}$ is nonzero on $X_{\lambda_{1}}$, with $s_{\ell}$ in its support. Moreover, as in [BE, Section 3] and the generalization appearing in [Ga2, Proposition 3.2], we know that if we zoom in on the measure $\mu^{r+1}_{\lambda_{1}}$ on $X_{\lambda_{1}}$ at the parameter $s_{\ell}$ at an exponential rate in $n$ (determined by the multipliers of the repelling cycles in the forward orbits of the marked critical points), and rescale the measure by $d^{nr}$, we get a nontrivial limit.
On the other hand, the critical point $c_{r+1}$ is attracted to a parabolic cycle for $f_{\ell}$ at the parameter $x_{1}$, by construction. By (4.8), we know that $\mu_{\lambda_{1}}^{1}=\alpha^{-1}\mu_{\lambda_{1}}^{r+1}\not=0$ on $X_{\lambda_{1}}$. But because of the parabolic point, we claim that if we zoom into the measure $\mu_{\lambda_{1}}^{1}$ on $X_{\lambda_{1}}$ at $s_{\ell}$ at the same exponential rate (which is supposed to give a nonzero limit), we actually get 0. This will produce our desired contradiction.
@@
Suppose $f_{x}$ is a holomorphic family of maps on $\mathbb{P}^{1}$ parametrized by $x\in\mathbb{D}^{r}\subset\mathbb{C}^{r}$, with a marked critical point $c:\mathbb{D}^{r}\to\mathbb{P}^{1}$. Assume that $c(0)$ is attracted to a parabolic fixed point at $z=0$ for $f_{0}$. Then for any constant $\rho>1$, there exists $\delta>0$ so that the sequence of functions

### v3

@@ 2.4. The critical height
defines a function on $\mathcal{M}_{d}^{cm}$, which Silverman called the critical height and conjectured that it is comparable with a Weil height on $\mathrm{M}_{d}\setminus\mathcal{L}_{d}$; see, for example, [Si2]. This comparability was proved by Ingram in [In].
@@
Yuan and Zhang use their general equidistribution theorem [YZ, Theorem 5.4.3] to give an alternative proof of the arithmetic equidistribution of the points in $\mathrm{PCF}_{d}$ to the bifurcation measure $T_{\mathrm{bif}}^{\wedge 2d-2}$ in [YZ, §6.3], a result which was first obtained by Gauthier in [Ga3] (and was proved for the polynomial moduli spaces in [FG1]).
@@ 2.5. The Chow variety and maximal variation
Suppose that $\pi:\mathcal{X}\to V$ is a family of $r$-dimensional varieties in $Y$, parametrized by $\lambda$ in an irreducible quasi-projective variety $V$ of dimension $\ell\geq 1$. Assume that the generic fiber of $\mathcal{X}\to V$ is geometrically irreducible. We say that $\mathcal{X}\to V$ is maximally varying if the natural map from $V$ to the Chow variety $\mathrm{Ch}(Y,r,D)$ has finite fibers, where we send each irreducible fiber $X_{\lambda}:=\pi^{-1}(\lambda)$ to the point of $\mathrm{Ch}(Y,r,D)$ representing the cycle $\overline{X}_{\lambda}]$ on $\overline{Y}$.
@@
Now suppose $\ell>1$ and that we know the result for maximally varying families with base dimension $\ell$. Note that $\dim\mathcal{X}=\ell+r$ and $\dim\mathcal{X}_{V}^{\ell}=\ell(r+1)$. Choose any smooth point $x_{0}$ of $\rho_{1}(\mathcal{X})$ which is also smooth for a subvariety $X_{\lambda_{0}}$ it lies in, and which is regular for the maps $\rho_{1}:\mathcal{X}\to\rho_{1}(\mathcal{X})$ and $\pi_{1}:\rho_{\ell}(\mathcal{X}_{V}^{\ell})\to\rho_{1}(\mathcal{X})$ where the latter is the restriction of the projection $Y^{\ell}\to Y$ to the first coordinate. Here, ‘regular for the maps’ means that the fiber over $x_{0}$ has the expected dimension and there is a point in the fiber over $x_{0}$ mapping surjectively on tangent spaces to that of $x_{0}$.
@@ 3. Bifurcation measures on $\mathcal{X}_{V}^{\ell}$ and $\mathcal{X}_{V}^{\ell+1}$
In this section we construct nontrivial measures on $\mathcal{X}_{V}^{\ell}$ and $\mathcal{X}_{V}^{\ell+1}$ that are pullbacks of powers of the bifurcation currents in $\mathcal{M}_{d}^{cm}$, defined in §2.3. The measure on $\mathcal{X}_{V}^{\ell}$ will be defined in terms of $T_{\mathrm{bif}}$ of (2.3), while two measures on $\mathcal{X}_{V}^{\ell+1}$ will additionally involve choices of $r$-tuples of marked critical points. The proofs build on the strong rigidity result that the power $T_{\mathrm{bif}}^{\wedge r}$ of the bifurcation current is nonzero on every $r$-dimensional subvariety of $\mathrm{M}_{d}\setminus\mathcal{L}_{d}$, where $\mathcal{L}_{d}$ is the flexible Lattès locus; this is proved in [GOV, Lemma 6.8] for all $r$. (See the discussion in §2.3.)
We also appeal to descriptions of the support of powers of $T_{\mathrm{bif}}$ in $\mathrm{M}_{d}$ in terms of critical orbit behavior obtained by Buff-Epstein in [BE] and Dujardin in [Du]. Precisely, suppose that $f_{1},\ldots,f_{m}$ is a collection of holomorphic families of maps on $\mathbb{P}^{1}$ parameterized by $\lambda$ in a complex manifold $\Lambda$, and let $\Phi:=(f_{1},\ldots,f_{m}):\Lambda\times(\mathbb{P}^{1})^{m}\to\Lambda\times(\mathbb{P}^{1})^{m}$ be the induced holomorphic map. We say that an $m$-tuple of marked points ${\bf a}=(a_{1},\ldots,a_{m}):\Lambda\to(\mathbb{P}^{1})^{m}$ is transversely (respectively, properly) prerepelling for $\Phi$ at $\lambda_{0}$ over $\Lambda$ if there exist repelling periodic points $p_{i}$ of $f_{\lambda_{0}}$ and integers $n_{i}\geq 1$ such that $f_{\lambda_{0}}^{n_{i}}(a_{i}(\lambda_{0}))=p_{i}$ for each $i=1,\ldots,m$, and the following holds: if $p_{i}:U\to\mathbb{P}^{1}$ denotes the holomorphic continuation of the repelling periodic point $p_{i}=p_{i}(\lambda_{0})$ over some neighborhood $U\subset\Lambda$ of $\lambda_{0}$ where it remains repelling, then the graphs of the maps
$\lambda\longmapsto\big(f_{1,\lambda}^{n_{1}}(a_{1}(\lambda)),\ldots,f_{m,\lambda}^{n_{m}}(a_{m}(\lambda))\big)\quad\text{and}\quad\lambda\longmapsto(p_{1}(\lambda),\ldots,p_{m}(\lambda))$
intersect transversely (respectively, properly) in $U\times(\mathbb{P}^{1})^{m}$ at $\lambda_{0}$.
For positive integers $m\leq 2d-2$, marked critical points ${\bf c}=(c_{1},\ldots,c_{m})$, and $f_{1}=\cdots=f_{m}:=f$, Dujardin proved that there is a dense set of parameters in $\operatorname{supp}T_{1}\wedge\cdots\wedge T_{m}$ in $\mathcal{M}_{d}^{cm}$ at which $\bf c$} is transversely prerepelling over $\mathcal{M}_{d}^{cm}$ [Du, Theorem 1.1]; recall that $T_{i}$ was defined in (2.2). And conversely, if the $m$-tuple of critical points is properly prerepelling for $(f,\ldots,f)$ at some parameter $\lambda_{0}$ over a parameter space of dimension $\geq m$, then $T_{1}\wedge\cdots\wedge T_{m}$ must be nonzero in that parameter space with $\lambda_{0}$ in its support, as first observed in [BE] (where it was proved in the case m=2d2m=2d-2 and assuming transversely prerepelling) and revisited in [Ga1, Theorem 6.2], [Ga2, Theorem 2.3], and [BBD, Proposition 3.7] in more general contexts and without assuming transversality. See [DM1, Proposition 4.8] for the version we use, which implies that the latter converse result remains true with any $m$-tuple of maps $(f_{1},\ldots,f_{m})$ of degree $d$ in place of $(f,\ldots,f)$.
@@ 3.1. Measure on $\mathcal{X}_{V}^{\ell}$
so that the measure $\mu_{\ell}\not=0$. Here we used the fact that $\dim\mathcal{X}_{V}^{\ell}=\dim\mathcal{Y}_{\ell}$ from the maximal variation.
@@ 3.2. Measures on $\mathcal{X}_{V}^{\ell+1}$
be as in Proposition 3.1 parameterized by $\mathcal{X}_{V}^{\ell}$ via the map $\rho_{\ell}$ of (3.1), and let
@@
be the family of maps on $(\mathbb{P}^{1})^{\ell(r+1)+r}$ parameterized by $\mathcal{X}_{V}^{\ell+1}$, where $f_{1},\ldots,f_{\ell}$ are each repeated $r+1$ times, but the final coordinate $f_{\ell+1}$ is only repeated $r$ times. Specializing at each $\lambda\in V$ we also have a family of maps
$F^{(\lambda)}:=(f,\ldots,f):X_{\lambda}\times(\mathbb{P}^{1})^{r}\to X_{\lambda}\times(\mathbb{P}^{1})^{r},$
parametrized by $X_{\lambda}\subset\mathcal{M}_{d}^{cm}$, and, focusing on the $\ell$-th factor of $\mathcal{X}_{V}^{\ell}$ and associated components of $\bf c$}, we will use the marked critical points
$c_{\ell,1}|_{X_{\lambda}},\ldots,c_{\ell,r+1}|_{X_{\lambda}}:X_{\lambda}\to\mathbb{P}^{1}$
to define two measures $\mu_{\ell}^{i}$ and $\mu_{\ell}^{j}$ on $\mathcal{X}_{V}^{\ell+1}$. Conditions (1) and (2) of Proposition 3.2 will allow us to conclude that $\mu_{\ell}^{i}$ and $\mu_{\ell}^{j}$ are nonzero. Conditions (3) and (4) will allow us to prove in Section 4 that $\mu_{\ell}^{i}$ and $\mu_{\ell}^{j}$ are not proportional measures.
@@
    the $r$-tuple $(c_{\ell,1},\ldots,\widehat{c_{\ell,i}},\ldots c_{\ell,r+1})$ is transversely prerepelling for $F^{(\lambda_{0})}$ at $t_{\ell}$ over $X_{\lambda_{0}}$;
@@
    the $r$-tuple $(c_{\ell,1},\ldots,\widehat{c_{\ell,j}},\ldots c_{\ell,r+1})$ is properly prerepelling for $F^{(\lambda_{0})}$ at $t_{\ell}$ over $X_{\lambda_{0}}$;
@@
    the $r$-tuple $(c_{\ell,1},\ldots,\widehat{c_{\ell,i}},\ldots c_{\ell,r+1})$ is transversely prerepelling for $F^{(\lambda_{1})}$ at $s_{\ell}$ over $X_{\lambda_{1}}$; and
  4. (4)
    the critical point $c_{\ell,i}$ is attracted to a parabolic periodic point of $f$ at $s_{\ell}\in X_{\lambda_{1}}$,
where $\widehat{c_{\ell,i}}$ means that the $i$-th coordinate is omitted. Let $\Pi_{1}:\mathcal{X}^{\ell}_{V}\times_{V}\mathcal{X}\to\mathcal{X}_{V}^{\ell}$ and $\Pi_{2}:\mathcal{X}_{V}^{\ell}\times_{V}\mathcal{X}\to\mathcal{X}$. Then
@@
By the transversality, we know that $r$ of the remaining critical points, say $c_{\ell,1},\ldots,c_{\ell,r}$, are transversely prerepelling at $t_{\ell}$ for $F^{(\lambda_{0})}=(f,f,\ldots,f)$ over $X_{\lambda_{0}}$.
We now aim to show that by modifying the parameter $x_{0}$ if necessary, we can arrange so that another $r$-tuple, say $c_{\ell,2},\ldots,c_{\ell,r+1}$, is also properly prerepelling at $t_{\ell}$ for $F^{(\lambda_{0})}$ over $X_{\lambda_{0}}$. Note that transversality of the intersection $S_{1,1}\cap\cdots\cap S_{\ell-1,r+1}$ at the original $x_{0}$ persists in a neighborhood, so the first $(\ell-1)(r+1)$ critical points remain transversely prerepelling in a neighborhood of $x_{0}$ for $(f_{1},\ldots,f_{\ell-1})$.
Consider the intersection
@@
We conclude that we can adjust our parameter $\lambda_{0}$ slightly, if needed, so that points $c_{\ell,1},\ldots,c_{\ell,r}$ are transversely prerepelling at a new choice of $x_{0}\in\Gamma$ along $X_{\lambda_{0}}$ while $c_{\ell,2},\ldots,c_{\ell,r+1}$ are properly prerepelling at $x_{0}$ along $X_{\lambda_{0}}$. This completes the proofs of (1) and (2) of the proposition.
As $c_{\ell,r+1}$ is bifurcating along the curve we called $\Gamma$, the characterizations of $J$-stability (as in [MSS]) imply that there are parameters, everywhere dense in its bifurcation locus, for which the maps have a parabolic cycle; the theory of renormalization (see [DH1]) implies that $c_{\ell,r+1}$ will lie in the basin of the parabolic cycle. We let $\lambda_{1}$ be a parameter close to $\lambda_{0}$ in the support of $\pi_{*}\mu_{\ell}$, where $\pi$ is the projection to $V$, so that $c_{\ell,r+1}$ is attracted to a parabolic cycle at $x_{1}:=\Gamma\cap X_{\lambda_{1}}$, while $c_{\ell,1},\ldots,c_{\ell,r}$ remain transversely prerepelling at $x_{1}$ along $X_{\lambda_{1}}$. This completes the proof of properties (1)–(4).
To prove the non-vanishing of $\mu_{\ell}^{i}$ and $\mu_{\ell}^{j}$, consider the map $\Phi_{\ell+1}=f_{1}^{\times(r+1)}\times\cdots\times f_{\ell}^{\times(r+1)}\times f_{\ell+1}^{\times r}$ on $(\mathbb{P}^{1})^{\ell(r+1)+r}$ over $\mathcal{X}_{V}^{\ell+1}$. The argument above, leading to properties (1) and (2), shows that we can label critical points of $f_{1},\ldots,f_{\ell+1}$ so that the graphs of the marked points
@@
Suppose that $\mathcal{X}\to V$ is a maximally varying family of $r$-dimensional varieties $X_{\lambda}\subset M_{d}$ parametrized by $V$ of dimension $\ell\geq 1$, defined over $\overline{\mathbb{Q}}$. As in §2.1, we may pass to a finite branched cover $\mathcal{M}_{d}^{cm}\to\mathrm{M}_{d}$ over which we can define a family of maps $f:\mathbb{P}^{1}\to\mathbb{P}^{1}$ with 2d22d-2 marked critical points. Pulling back each $X_{\lambda}$ and possibly replacing $V$ with a finite branched cover, we will continue to use the notation $\mathcal{X}\to V$ for the family of varieties in $\mathcal{M}_{d}^{cm}$. Therefore, we may assume that we can define the family of product maps
@@
In what follows it is convenient to work with the Deligne pairing (of relative dimension dimension 0), of the invariant metrized line bundles restricted to graphs of critical marked points (see [YZ, §4.2] for metrics of the Deligne pairing and [YZ, §6.2.2] for the ‘restriction’ of the invariant metrized line bundle to a subvariety). By [YZ, Theorem 4.2.3], to each marked tuple of critical points ${\bf c}:\mathcal{X}^{\ell}_{V}\to(\mathbb{P}^{1})^{\dim\mathcal{X}^{\ell}_{V}}$ we can associate a metrized line bundle $\langle\overline{L}_{\Phi}|_{\Gamma_{\bf c}}\rangle\in\widehat{\mathrm{Pic}}(\mathcal{X}^{\ell}_{V})_{\mathbb{Q},\mathrm{nef}}$, where $\Gamma_{\bf c}$ denotes the graph of $\bf c$} in $\mathcal{X}^{\ell}_{V}\times(\mathbb{P}^{1})^{\dim\mathcal{X}^{\ell}_{V}}$. Moreover, we let
@@
on $\mathcal{X}^{\ell+1}_{V}$ for $i=1,2$. By [YZ, Lemma 6.2.1, Theorem 4.1.3], the heights associated to $\overline{M}_{i}$ vanish at PCF points; see also [YZ, Lemma 6.3.2]. More precisely, for a point $(\lambda_{0},t_{1},\ldots,t_{\ell+1})\in\mathcal{X}^{\ell+1}_{V}(\overline{\mathbb{Q}})$ we have
@@
(4.7) $\int_{\mathcal{X}^{\ell}_{V}}\left(\int_{X_{\pi(x)}}\phi\cdot\bigwedge_{i=1}^{r}T_{f_{\ell+1},c_{i}}|_{X_{\pi(x)}}\right)d\mu_{\ell}(x)=\alpha\int_{\mathcal{X}^{\ell}_{V}}\left(\int_{X_{\pi(x)}}\phi\cdot\bigwedge_{i=2}^{r+1}T_{f_{\ell+1},c_{i}}|_{X_{\pi(x)}}\right)d\mu_{\ell}(x)$
@@
as measures on $X_{\lambda}$, for some constant $\alpha>0$ and for all $\lambda$ in the support of $\pi_{*}\mu_{\ell}$ in $V$. Indeed, suppose there exists $b\in V(\mathbb{C})$ where (4.8) doesn’t hold, and so that $\pi_{*}\mu_{\ell}(U)>0$ for every open neighborhood $U$ of $b$. Then we can find a continuous, non-negative, compactly supported function $\psi$ on $\mathcal{X}$ such that
$\int_{X_{b}}\psi\,d\mu^{r+1}_{b}(x)\neq\alpha\int_{X_{b}}\psi\,d\mu^{1}_{b}(x).$
@@
The remainder of the proof is devoted to deriving a contradiction from the equality (4.8). We first sketch the idea. Recall from Proposition 3.2 that we have parameter $x_{1}=(\lambda_{1},s_{1},\ldots,s_{\ell})\in\operatorname{supp}\mu_{\ell}$ so that the tuple of marked critical points $(c_{1},\ldots,c_{r})$ is transversely prerepelling for $F^{(\lambda_{1})}=(f,\ldots,f)$ at $s_{\ell}$ over $X_{\lambda_{1}}$. This implies that $\mu_{\lambda_{1}}^{r+1}$ is nonzero on $X_{\lambda_{1}}$, with $s_{\ell}$ in its support by [BE]. Moreover, as in [BE, Section 3] and the generalization appearing in [Ga2, Proposition 3.2], we know that if we zoom in on the measure $\mu^{r+1}_{\lambda_{1}}$ on $X_{\lambda_{1}}$ at the parameter $s_{\ell}$ at an exponential rate in $n$ (determined by the multipliers of the repelling cycles in the forward orbits of the marked critical points), and rescale the measure appropriately we get a nontrivial limit; see (4.11). On the other hand, we know from Proposition 3.2 that the critical point $c_{r+1}$ is attracted to a parabolic cycle at $s_{\ell}\in X_{\lambda_{1}}$. Zooming into the measure $\mu_{\lambda_{1}}^{1}$ on $X_{\lambda_{1}}$ at $s_{\ell}$ at the same exponential rate (which is supposed to give a nonzero limit), we actually get 0 from the following lemma. This will produce our desired contradiction.
@@
Suppose that $f:\mathbb{D}^{r}\times\mathbb{P}^{1}\to\mathbb{D}^{r}\times\mathbb{P}^{1}$ is a holomorphic family of maps with degree $d\geq 2$ parametrized by $\mathbb{D}^{r}\subset\mathbb{C}^{r}$ with a marked critical point $c:\mathbb{D}^{r}\to\mathbb{P}^{1}$. Assume that $c(0)$ is attracted to a parabolic fixed point at $z=0$ for $f_{0}$. Then for any constant $\rho>1$, there exists $\delta>0$ so that the sequence of functions